# Hydra: Virtualized Multi-Language Runtime for High-Density Serverless Platforms

### v2

@@ 4.1. Node Manager
Instead of creating a Hydra instance on-demand right before serving an incoming invocation, the node manager fetches a Hydra instance from a fixed configurable pool of pre-allocated Hydra instances. The node manager is responsible for maintaining this pool. A periodic background job monitors the state of this pool to prevent the pool from complete depletion. If the pool load reaches 30% of the initial capacity, this job reclaims 15% of least-recently-used warm Hydra instances to replenish the pool with fresh instances. The rest of the warm Hydra instances remain untouched and ready to serve future invocations. Such an approach reduces the chances of experiencing a cold start on a critical path of a function invocation, thus reducing the tail latency of the function requests.
4.2. Scheduler
The scheduler governs a fixed number of nodes and keeps track of the state of each node to make appropriate scheduling decisions and balance the load. The state of a node includes up-to-date information about the node’s active invocations and registered tenants and functions. The scheduler is aware of the Hydra’s co-location capabilities. Instead of merely balancing the load, it tries to group functions coming from the same tenant in the same nodes to make the best use of Hydra at the node level.
The scheduler estimates the memory utilization of each node and ensures that by sending a new invocation into a node, it will not exceed the node’s memory limit. The platform supports three main runtime modes: tenant-based co-location (default for Hydra), single-function co-location (only invocations of the same function can be co-located), and no co-location. For each mode, memory utilization is calculated in a different way. Due to space constraints, we focus on tenant-based co-location mode, which can be calculated as follows:
$mem_{node}=(\sum_{t}^{T}\lceil\sum_{f}^{F}(size_{f}\times inv_{f})/size_{i}\rceil)\times size_{i}$
where
$mem_{node}$  : memory consumption of the node
$size_{f}$  : maximum memory for the function f
$inv_{f}$  : number of active invocations of the function f
$size_{i}$  : maximum memory footprint of the instance
First, the overall memory consumption of active functions for each tenant t is calculated. Then, this value is divided by the size of the Hydra instance and rounded up to get the number of Hydra instances for each tenant t. Finally, the total number of all tenants’ instances is multiplied by the size of the instance to get the node’s memory consumption.
@@ 5. Evaluation
This section studies the performance of Hydra and compares it to other state-of-the-art runtimes. We start by measuring the density of different runtimes using a variety of serverless benchmarks (section 5.1). Then, we replay a real serverless invocation trace on a local platform (section 5.2) to assess the platform-wide impact of Hydra and our tenant-based scheduling algorithm. Table 1 lists the benchmarks used to evaluate Hydra. We combine benchmarks from Sebs (copik:2021, ) (Python and JavaScript benchmarks) and Photons (dukic:2020, ) (Java benchmarks).
@@
JavaScript (js) helloworld (hw), dynamic-html (html), upload, thumbnail
@@ 5.1. Function Sandbox Density
Figure 8. Invocation latency and memory footprint comparing a new sandbox (Cold) to a restored snapshot (Snap).
@@
This experiment analyzes the density of four different types of serverless runtimes:
@@
  • OpenWhisk (openwhisk, ) runtimes used to handle Java (JVM), JavaScript (NodeJS), and Python (CPython) invocations. These runtimes are state-of-the-art, native language implementations. Each of these runtimes can handle invocations of a single function, one invocation at a time;
@@
  • NIFT, a variant of Hydra that does not allow multiple functions to execute in the same instance. This baseline mimics what is possible to achieve today by combining GraalVM Native Image and Truffle. Similarly to OpenWhisk runtimes, this variant can handle invocations of a single function, one invocation at a time;
@@
  • Forking, a variant of Hydra inspired by SAND (akkus:2018, ). In this system, multiple invocations of a single function can be served concurrently. Each invocation will be handled in a fork of the parent process that pre-loaded the function code. We use this runtime to compare the overheads of process forking compared to isolate sandboxes;
@@
  • Hydra, capable of handling multiple function invocations from different functions at the same time. We limit Hydra to use isolate memory sandboxes.
@@
Our goal is to measure how much throughput can be produced on a single core with 2 GBs of memory (we use a 1:2 core-to-memory ratio inspired by the ratio used in Amazon Lambda (lambda, )). Our target metric, density, is defined as the ratio of throughput per memory (ops/GB-sec). Since several benchmarks are I/O-bound (e.g., jv/hash, js/uploader, py/hw), we allow concurrent requests to share those resources to improve efficiency. To do so, we add additional concurrent requests until the efficiency of the runtime stops increasing. For OpenWhisk and NIFT, handling multiple requests concurrently means having several instances running with a fraction of a core and a fraction of 2 GBs. For Hydra, and Forking, we use a single instance with multiple sandboxes.
@@
Experiments are conducted on a single cluster node running Linux kernel 5.15.0 equipped with an Intel(R) Xeon(R) Gold 532 and 188GB of DDR4 DRAM. CPU frequency scaling, hyper-threading are disabled, and runtimes are pinned to a single core. Each runtime runs inside a Docker container with the set limits in terms of CPU and memory. Throughput is measured using Apache Bench, [^2] https://httpd.apache.org/docs/2.4/programs/ab.html and memory represents sum of the RSS of the runtime (and children processes if any). Results are presented in Figure 7. Each bar represents an average of 5 iterations where each iteration includes 25 requests.
@@ 5.1.1. Java Benchmarks
Hydra presents higher efficiency compared to other runtimes. Compared to OpenWhisk, for Java IO-bound benchmarks (hw, hash, and rest), Hydra reduces the number of runtime instances to one, therefore significantly reducing the memory footprint. In addition, CPU multiplexing improves by having a full core shared across several isolates compared to having multiple quotas of a core for each container. Forking has the closest performance compared to Hydra. The slight density degradation comes mostly from frequent context switching and the increased memory footprint that comes from duplicated pages in the child process.
Figure 9. Replaying a serverless invocation trace from Azure on Hydra with different scheduling strategies. The data is aggregated across all nodes.
@@ 5.1.2. JavaScript and Python Benchmarks
Similarly to Java benchmarks, Hydra is a system with higher throughput and lower memory footprint in most benchmarks. It allows multiple isolates to share the same Truffle JIT compiler and benefits from sandbox snapshotting to bypass JIT-compilation overheads. Hydra is followed by Forking and by NIFT. In several Python benchmarks, however, OpenWhisk registers higher density. After analyzing these runs we conclude that these results stem from some performance problems from Truffle’s Python implementation that is still in experimental mode. These problems lead mainly to low throughput and high memory footprint in particular in py/html, py/thumbnail, py/compress, and py/video. Overall, Hydra improves the average density by 4.47$\times$ compared OpenWhisk runtimes.
@@ 5.1.3. Function Snapshotting Impact on Density
We now evaluate the impact of function snapshotting in Hydra by comparing the latency and memory footprint required to handle a cold start (launching a new sandbox and invoking the function). Note that for sandboxes created from scratch, this may require code interpretation and JIT-compilation.
Results are depicted in Figure 8. Bot plots include Cold (vanilla sandboxes that suffer from a cold start), and Snap (restored sandboxes that restore the execution past a cold start). The latency in Snap includes both the restore latency and the function request latency. We omit some benchmarks (such as jv/video, jv/classify) that require process forking, something that our snapshotting mechanism does not support yet. Also note the log-scale on both y-axis. Overall, snapshotting reduces the latency to serve the first request by 36.62$\times$ and reduces the memory footprint by 3.15$\times$.
5.2. Real-World Experiment
So far, we have analyzed the performance of different serverless runtimes when executing a single function. However, in real environments, platforms deal with concurrent requests from different functions. To realize such an environment, we take advantage of the public Azure Functions trace (sharad:2020, ). The trace provides data about function invocations, including the function identifier, the duration, the memory footprint, and the tenant associated with the function.
@@ 5.2.1. Top-Level Scheduling Setup
Figure 11. Replaying a serverless invocation trace from Azure on different serverless runtimes.
As stated in section 4.2, the top-level component of the platform schedules requests in an attempt to group functions from the same tenant in the same node. To show that grouping tenants is more beneficial when using Hydra compared to a traditional round-robin scheduler, we simulated the execution of the Azure trace for 10 minutes with 200 nodes; each node has the capacity for running concurrent 40 Hydra instances, and each instance can co-locate eight parallel invocations. Figure 9 depicts the difference between the Hydra-aware and round-robin scheduling with regard to the number of cold starts and the memory consumption by the active invocations across all nodes. We can observe that the number of cold starts has decreased by 2.8 times, and the memory footprint of active invocations has decreased by 3.6 times on average.
According to the Azure Functions study (sharad:2020, ), a minority of functions generates a vast majority of all invocations. Therefore, with the tenant-grouping scheduling, the distribution of invocations and registered functions across the nodes of the serverless platform may be non-uniform, as some functions generate more load than others. Figure 10 demonstrates that the nodes with the smaller indexes tend to have few functions registered, albeit these functions get more invocations than others. The workers with the bigger indexes observe the opposite situation: they tend to have more different functions registered, but these functions are invoked infrequently.
@@ 5.2.2. Reproducing a Real-World Trace
We select a 1-hour window that best represents the average user activity in serverless. We reproduce this segment on a local cluster machine running Ubuntu 18.04.6 LTS (Linux kernel 5.8.5-050805) equipped with an Intel(R) Xeon(R) Silver 4114 CPU @ 2.20GHz and 128GB of DDR4 DRAM. CPU frequency scaling and hyper-threading are disabled. As the trace does not describe the actual function code executed, we use a Java-based function that performs hashing of the input data.
In this experiment, we replay the same serverless invocation trace with four execution modes: Hydra (co-locating all functions from the same tenant), Hydra single-function (co-locating only parallel invocations of the same function), Hydra single-invocation (not co-locating invocations), and the OpenWhisk (openwhisk, ) runtime. With the Hydra and Hydra single-function modes, each node of the serverless platform runs up to 40 Hydra instances, each instance can co-locate up to 8 concurrent isolated invocations, and each instance uses up to one core and two GBs of memory. With the Hydra single-invocation and OpenWhisk modes, each node runs up to 320 Hydra or OpenWhisk instances respectively, and each instance uses up to 1/8 of a core and 256 MBs of memory.
@@
Results in Figure 11 include the number of active instances (top left), the total memory consumption of the node (bottom left), the number of hot (picking a previously used instance), warm (fetching an instance from the pool described in section 4.1), and cold (launching a new instance) starts of the instances (top right), and the user request latency (bottom right). The left part of each plot shows the results of the trace execution from the node that has few frequently invoked functions (Node #1), and the right part shows the results from the node with hundreds of infrequently invoked functions (Node #2).
Overall, by being able to consolidate invocations of different functions in Hydra instances, it significantly reduces the number of active instances without degrading throughput, leading to memory reduction of up to 2.1 times on average compared to OpenWhisk in both nodes. In Node #2, Hydra single-function mode requires more instances than plain Hydra since the latter mode offers wider opportunities for co-location. The memory footprint of the Hydra single-invocation mode is comparable to the OpenWhisk execution as these modes use the same number of instances, but Hydra instances have a smaller footprint.
In Node #1, plain Hydra and Hydra single-function have similar opportunities to co-locate as there are only a few popular functions. Thus, these modes experience fewer cold starts compared to the execution modes that do not support co-location. However, since Node #2 has hundreds of infrequently invoked functions registered, the 40 Hydra instances are not fully utilized and are constantly re-created to accommodate all functions and tenants. The lower number of launched stacks results in reduced tail latency. Compared to OpenWhisk, Hydra reduced the number of cold starts by 48 times in Node #1 and by four times in Node #2. Figure 11 shows that user request latency is reduced by Hydra in both nodes.
5.3. Discussion
Hydra presents key advantages compared to other runtimes. First, it deploys AOT-compiled function code, which is faster to boot, has no warmup time, and requires less memory footprint (section 5.1.1). It also allows Truffle languages to reuse profiles and optimize code across invocations, further reducing redundancy to profile and optimize code (section 5.1.2). Finally, it can host multiple invocations concurrently, ensuring that the virtualization infrastructure is not duplicated, resulting in lower memory utilization and fewer cold starts (section 5.2).
@@ 6. Related Work
Runtime Forking and Snapshotting have been widely studied in the context of Serverless. SAND (oakes:2018, ), SOCK (akkus:2018, ) rely on process forking to minimize cold start invocation latency and reduce memory footprint by sharing runtime libraries. Other systems such as FaaSnap (ao:2022, ), SEUSS (cadden:2020, ), Catalyzer (du:2020, ), vHive (ustiugov:2021, ), Prebaking (silva:2020, ), Fireworks (shin:2022, ), and Medes (saxena:2022, ) proposed different snapshotting techniques to reduce cold start latency and memory footprint.
Hydra aggregates multiple function invocations inside lightweight sandboxes in a single virtualization stack. Hydra allows applications to use process-level isolation, but can also support in-process isolate sandboxes which are faster to launch, use less memory, and deliver higher throughput (section 5.1). Snapshotting is orthogonal to Hydra’s contributions and it can be used together with Hydra (as we do in section 5.1) to reduce the cold start latency of a new Hydra runtime. Snapshotting alone, however, does not reduce resource redundancy as multiple virtualization stacks are still launched from a common snapshot.
@@
Runtime Pre-warmup has also been studied with the goal of launching runtimes just before invocations are received (mahgoub:2022, ) and thus hiding the cold start latency. Others have studied how to reduce the pre-warmup cost by taking advantage of heterogeneous resources (roy:2022, ). Mohan et al. (mohan:2019, ) proposed a pool of containers that could be utilized when function invocations arrive. Runtime pre-warmup is yet another technique that could be used in combination with Hydra.
@@
Replacing multiple runtimes that implement a single language with a single runtime that implements multiple languages makes it easier to manage runtime images. By doing so, predicting which runtime to use for the next invocation becomes a lesser problem. In addition, a pool of Hydra runtimes could be used to accommodate any function invocation as long there is a Truffle implementation for the language the function is implemented in.
@@
Virtualizing Runtimes allows one or more functions to share a single runtime. Sharing, however, can be achieved with different levels of isolation. Crucial (pons:2019, ) and Nightcore (jia:2021, ) handle multiple function invocations in a single runtime using different threads. Photons (dukic:2020, ) also runs concurrent function invocations in a single runtime but applies Java bytecode transformations to ensure isolation (i.e., concurrent function invocations could not interact with other co-executing invocations). Boucher et al. (boucher:2018, ) rely on Rust’s memory safety guarantees to handle multiple invocations concurrently in isolation. Cloudfare Workers (cloudfare, ) rely on V8 memory isolates to run concurrent JavaScript functions in the same V8 runtime instance. Faasm (shillaker:2020, ) and Gackstatter et al. (gackstatter:2022, ) proposed using WebAssembly (haas:2017, ) to execute multiple invocations in isolation. WebAssembly however, still presents several limitations that prevent modern high-level languages (Java, JavaScript, and Python) from being supported (see section 2.2 for more details). As a result, runtimes such as Faasm need to compile a Python runtime such as CPython to WebAssembly in order to run Python functions. If concurrent invocations are to be handled, multiple runtimes will run concurrently in different WebAssembly execution environments leading to resource redundancy.
@@
Hydra enforces isolation through memory isolates. However, compared to Cloudfare Workers, Hydra is not restricted to JavaScript functions, it supports multi-thread function code, [^3] Multi-threading was disabled as a security measure resulting from designing the system to accommodate multiple tenants (workerd, ). it proposes AOT-compilation for user functions, and offers built-in process-level isolation. Opposed to Faasm, Hydra relies on Truffle to support multiple language support and therefore, concurrent invocations always share the underlying runtime. In addition, to further share the JIT compilation and profiling, Hydra shares the code caches among concurrent invocations of the same function.
Minimal Runtimes Although we use Firecracker VMs to isolate different tenants, the proposed techniques are not specific to Firecracker nor to system-VMs. Thus, container-based solutions such as Cntrl (thalheim:2018, ) and gVisor (gvisor, ) could also be used. Unikernel VMs (manco:2017, ; koller:2017, ; zhang:2018, ) are also a possible alternative to reduce Hydra’s cold start. However, supporting functions whose dependencies are only known at run-time creates additional challenges to reduce the components to include at VM image generation time. As a consequence, Hydra includes a full Linux kernel and requires functions to package any user-level library dependencies along with the function code.
@@
Profile and Code Cache Sharing reduce code profiling and compilation overheads. Profile sharing has been proposed by Arnold et al. (arnold:2005, ) and Ottoni et al. (ottoni:2021, ) as a way to re-use profiling information. By doing so, future invocations do not need to repeat all the code profiling steps and can benefit from the profiles gathered by previous invocations. A similar technique has also been used to share Just-in-Time compiled code (khrabrov:2022, ; xu:2018, ) among different invocations of the same application/function. Such techniques could be integrated into Hydra to share Truffle code caches. However, solutions such as JITServer (khrabrov:2022, ), ShareJIT (xu:2018, ) are designed to optimize long-running applications and are not prepared to deal with the challenges inherent to fast-executing and lightweight Serverless functions (carreira:2021, ).

### v3

@@ 4.2. Hydra Manager
A periodic background job monitors the state of the pool to prevent it from complete depletion. If the pool is depleted to a configurable threshold during execution, this job reclaims a configurable number of least-recently-used Hydra instances to replenish the pool with fresh instances. The rest of the warm Hydra instances remain untouched and ready to serve future invocations.
@@ 5. Evaluation
This section analyzes the performance of Hydra and compares it to other state-of-the-art runtimes and platforms. Hydra is implemented in a combination of Java and C in 10298 lines of code which we plan to make available to public. We start by measuring the density of different runtimes using a variety of serverless benchmarks (section 5.1). Then, we replay a real serverless invocation trace on a local Hydra platform deployment (section 5.2) to evaluate Hydra at the platform level. Table 1 lists the benchmarks used in our experiments. We combine benchmarks from Sebs (copik:2021, ) (Python and JavaScript benchmarks) and Photons (dukic:2020, ) (Java benchmarks).
@@
JavaScript (js) helloworld (hw), dynamic-html (html), upload
@@ 5.1. Function Sandbox Density
Figure 9. Invocation latency and memory footprint comparing a new sandbox (Cold) to a restored from snapshot (Snap) and warmed-up (Warm) in Hydra.
@@
This experiment analyzes the function sandbox density of four different types of serverless runtimes:
@@
  • OpenWhisk (openwhisk, ) runtimes used to handle Java (JVM), JavaScript (Node.js), and Python (CPython) invocations. These runtimes are state-of-the-art, native language implementations. Each of these runtimes can handle invocations of a single function, one invocation at a time;
  • Knative (knative, ) runtimes are similar to OpenWhisk but capable of handling multiple concurrent invocations of the same function. Parallel invocations are not isolated from each other. Another difference is that it uses GraalVM Native Image (graalvm, ) to compile Java functions into native executables. For Python and JavaScript, Knative uses CPython and Node.js by default, respectively;
@@
  • NITF, a variant of Hydra that does not allow multiple functions to execute in the same instance. This baseline mimics what is possible to achieve today by combining GraalVM Native Image and Truffle. Similarly to OpenWhisk runtimes, this variant can handle invocations of a single function, one invocation at a time;
@@
  • Forking, a variant of Hydra inspired by SAND (akkus:2018, ). In this system, multiple invocations of a single function can be served concurrently. Each invocation will be handled in a fork of the parent process that pre-loaded the function code. We use this runtime to compare the overheads of process forking compared to sandboxes;
@@
  • Hydra, capable of handling multiple function invocations from different functions at the same time. We allow Hydra to use already snapshotted sandboxes (see Figure 9 where we measure the effect of snapshotting).
@@
Our goal is to measure how much throughput can be produced on a single core with 2 GBs of memory  [^2] The used ratio is inspired by the 1:769 core-to-memory fixed ratio used in Amazon Lambda (lambdaquotas, ), which we approximate to 1:2 for simplicity.. Our target metric, density, is defined as the ratio of throughput per memory (ops/GB-sec). Naturally, higher density can be achieved through higher throughput or lower memory consumption. Since several benchmarks are I/O-bound (e.g., jv/hash, js/uploader, py/hw), we allow concurrent requests to share the virtualization stack to improve density. To do so, we add additional concurrent requests until the density of the runtime stops increasing. For OpenWhisk and NITF, handling multiple requests concurrently means having several instances running with a fraction of a core and a fraction of 2 GBs. For Hydra and Forking, we use a single instance with multiple sandboxes.
@@
Experiments are conducted on a single cluster node running Linux kernel 5.15.0 equipped with 2x Intel Xeon Gold 5320 CPU @ 2.2GHz and 128GB of DDR4 DRAM. CPU frequency scaling and hyper-threading are disabled, and runtimes are pinned to a single core. Each runtime runs inside a Docker container with limited CPU and memory. Throughput is measured using Apache Bench, [^3] https://httpd.apache.org/docs/2.4/programs/ab.html and memory represents sum of the RSS of the runtime (and children processes if any). The results are presented in Figure 8. Each bar represents an average of five iterations, where each iteration includes 10-16000 requests depending on the duration of each function invocation (which ranges from a few milliseconds to a few seconds).
@@ 5.1.1. Java Benchmarks
[…]
@@ 5.1.2. JavaScript and Python Benchmarks
[…]
@@ 5.1.3. Function Snapshotting Impact on Density
We now evaluate the impact of sandbox snapshotting in Hydra by comparing the latency and memory footprint required to handle a cold start (launching a new sandbox and invoking the function). Note that for sandboxes created from scratch, this may require code interpretation and JIT-compilation.
[…]
5.1.4. Function Checkpointing Overhead
[…]
5.2. Reproducing a Serverless Trace
So far, we have analyzed the performance of different serverless runtimes when executing multiple invocations a single function. However, in real environments, platforms deal with concurrent requests from different functions. To implement such an environment, we take advantage of the public Azure Functions trace (sharad:2020, ). The trace provides data about function invocations, including the function identifier, duration, memory footprint, and the tenant associated with the function.
@@ 5.2.1. Top-Level Scheduling Setup
Figure 10. Serverless traces, downscaled with random or statistic-driven approaches to 25% of the original trace, will only use 25% of the platform’s control plane, whereas running the full trace will use 100% of the platform’s control plane. Blue color represents actively used capacity, and gray represents unused capacity.
@@
Existing approaches to evaluating the impact of virtualization stacks in a realistic serverless scenario often involve random or statistic-driven down-sampling (ustiugov:2023, ; zhang:2023, ; singhvi:2021, ; miao:2024, ; fuerst:2021, ) of the real-world serverless traces to run workloads on a lower scale. While these down-sampling techniques allow pushing the load to the desired portion of the serverless platform’s data plane (i.e., invocations served in a particular node or worker), the load on the control plane (i.e., invocation scheduling) is partially omitted (see Figure 10, left side). Therefore, we argue that such approaches are applicable at the node or worker level but do not account for the platform scheduling that, depending on its policies, may impact the way virtualization stacks are utilized. Besides, random or statistic-driven down-sampling may require multiple experimental iterations to select a sample trace for a particular testbed, as they do not adapt with back-pressure when the node struggles with a particular set of requests.
The scheduler described in section 4.1, in addition to running workloads in production, can also be used as a benchmarking tool for the serverless platform. In an experimental setup, the scheduler runs the original (full) trace. The nodes (some portion or even all of them) can simulate serving the invocations instead of running the actual functions, thus not requiring computational resources. This way, the control plane is loaded realistically since the scheduler processes the entire original trace (see Figure 10, right side). Besides, the scheduler adapts to back-pressure that prevents it from overloading nodes since the scheduler keeps track of the state of each node.
Analyzing the Azure Function trace (sharad:2020, ), it is possible to demonstrate that a minority of functions generates the vast majority of all invocations. Therefore, with colocation-oriented scheduling, the distribution of invocations and registered functions across the nodes of the serverless platform can be nonuniform, as some functions generate more load than others. Figure 11 demonstrates that some nodes tend to have few functions registered, albeit these functions get more invocations than others. Other nodes observe the opposite situation: they tend to have more different functions registered, but these functions are invoked infrequently. With this insight, we sample the performance of the Hydra platform in two different nodes, one with low heterogeneity (invocations dominated by a few functions, labeled Node #1) and another with high heterogeneity (invocations originate from many different functions, labeled Node #2).
(a) Number of active runtimes.
(b) System memory footprint.
(c) End-to-end user request latency.
Figure 12. Replaying a serverless invocation trace from Azure (sharad:2020, ) on different serverless runtimes. Node #1 served 167043 invocations of 157 functions from 86 tenants. Node #2 served 23267 invocations of 793 functions from 311 tenants.
@@ 5.2.2. Reproducing a Real-World Trace
We select a 30-minute window that represents the average invocation activity in serverless. We analyzed all other 30-minute windows throughout the same day of the trace to ensure that the selected trace exhibits the same common profile in terms of the number of invocations processed, and total functions and tenants registered. We reproduce the selected segment on a local cluster machine with the same specifications as in section 5.1. The scheduler runs on another machine with identical specifications. As the function workload, we use the benchmarks from Table 1. Each benchmark is configured to use limited resources according to its complexity, 1/8 CPU and 256 MB of memory being the minimal resource allocation for the function, and 1 CPU and 2048 MB being the maximum resource allocation. We assign languages to functions from the trace according to the recent study on runtimes in serverless (newrelic, ) to maintain a realistic language distribution, and assign benchmarks (workloads) to functions based on the average duration of each function from the trace.
In this experiment, we replay the same serverless invocation trace from the scheduler machine with four execution modes: Hydra* (colocating all invocations), Hydra (colocating invocations to functions of the same tenant), Knative (knative, ) (colocating invocations of the same function), and OpenWhisk (no colocation). The scheduler operates with 99 nodes that simulate invocations and one node that executes the actual function code. The Hydra platform maintains a pool of 416 pre-created Hydra workers (containers running Hydra runtimes) as described in section 4.2. The total memory footprint of this pool is 1.9 GB. For Knative and OpenWhisk, workers are created on demand since for OpenWhisk, the pool would have to be partitioned across languages, and the platform cannot resize workers (language runtimes) after they are instantiated, and for Knative, workers are function-specific (as discussed in section 2.3). For all modes, the platform does not run more than 416 concurrent workers and reclaims the least recently used ones when it needs to create more. The platform keeps each worker alive for a randomized time period between 10 and 20 minutes after the last invocation.
Mode Cold Warm
OpenWhisk 2821 0
Knative 209 0
Hydra 0 118
Hydra* 0 1
(a) Node #1 (167K invocations).
Cold Warm
3049 0
2064 0
0 408
0 4
(b) Node #2 (23K invocations).
Table 2. Cold and warm starts from replaying the serverless invocation trace.
@@
The results in Figure 12 include the number of active instances (a), the total memory consumption of the node (b), and the user request latency (c). In addition, Table 2 shows the number of warm (fetching a worker from the pool) and cold (launching a new worker) starts. The ”Node #1” label denotes the node serving a few frequently invoked functions, and the ”Node #2” label denotes the node with many infrequently invoked functions.
Overall, by being able to consolidate invocations of different functions in Hydra instances, it significantly reduces the number of active instances without degrading latency, leading to memory reduction of up to 21.3-40.4% on average compared to OpenWhisk in both nodes if using Hydra, and up to 35.6-43.9% if using Hydra*. In Node #1, Knative shows lower memory footprint compared to Hydra modes because only a few functions generate most of the load, and only one Knative worker is needed for each function. However, it is important to note that the lower memory footprint offered by Knative comes as a consequence of the lack of isolation between concurrent requests within each worker, meaning that the heap and the other runtime components will be shared. Hydra, on the other hand, runs concurrent invocations in sandboxes that do not share runtime components, leading to a higher memory footprint. In Node #2, Hydra reduces the memory footprint by 14.5%, and Hydra* reduces the memory footprint by 30% compared to Knative because Knative needs one worker per function, whereas Hydra colocates functions from the same tenant, and Hydra* colocates all functions.
Hydra reduces tail latency by 51.4-375.5 $\times$× (p99) in Node #2 compared to Knative and OpenWhisk, respectively (Figure 12.c), as it fetches pre-created Hydra workers from the pool instead of initializing them on demand. OpenWhisk and Knative, on the other hand, experience numerous cold starts (see Table 2.b) for many different functions invoked infrequently. In Node #1, where a few functions receive the majority of invocations, the impact of cold starts for Knative and, to a lesser extent, for OpenWhisk is amortized by the high number of invocations to the few functions that can reuse workers. Still, Hydra reduces p99 latency by 45.3 $\times$× compared to OpenWhisk since OpenWhisk needs to initialize and run more workers to accommodate concurrent requests, and by 1.9 $\times$× compared to Knative since Knative cannot colocate different functions even from the same tenant.
Summary. By combining a virtualized runtime, colocation-aware scheduler, and a local pool of pre-created generic runtime instances, Hydra significantly increases density, reduces memory footprint, and eliminates cold starts, thus reducing tail latency.
@@ 6. Related Work
Runtime Forking and Snapshotting have been widely studied in the context of serverless. SAND (oakes:2018, ) and SOCK (akkus:2018, ) rely on process forking to minimize cold start invocation latency and reduce memory footprint by sharing runtime libraries. Other systems such as FaaSnap (ao:2022, ), SEUSS (cadden:2020, ), Catalyzer (du:2020, ), vHive (ustiugov:2021, ), Prebaking (silva:2020, ), Fireworks (shin:2022, ), Medes (saxena:2022, ), and Sabre (lazarev:2024, ) proposed different snapshotting techniques to reduce cold start latency and memory footprint. Snapshotting alone, however, does not reduce resource redundancy as multiple virtualization stacks are still launched from a common snapshot.
Hydra aggregates multiple function invocations in lightweight sandboxes in a single virtualization stack. Doing so allows applications to use process-level isolation but can also support in-process memory sandboxes, which are faster to launch and lead to higher density (section 5.1). Hydra can additionally snapshot individual sandboxes (section 3.4) to reduce the cold start latency of a new Hydra function sandbox. Deciding when to checkpoint a particular sandbox is outside the scope of this work and solutions proposed in previous works (such as Pronghorn (kohli:2024, )) can be utilized.
@@
Runtime Pre-Warmup has also been studied with the goal of launching runtimes just before invocations are received (mahgoub:2022, ) and thus hiding the cold start latency. Others have studied how to reduce the pre-warmup cost by taking advantage of heterogeneous resources (roy:2022, ). Mohan et al. (mohan:2019, ) proposed a pool of containers that could be utilized when function invocations arrive. TrEnv (huang:2024, ) offered a pool of repurposable container resources that could be used with any function. Our platform pre-warmups entire Hydra instances (container and runtime) in the pool as described in section 4.2 but could also be integrated with existing works that predict future function invocations.
@@
However, replacing multiple runtimes that implement a single language with a single runtime that implements multiple languages also makes it easier to manage runtime images. By doing so, predicting which runtime to use for the next invocation becomes a lesser problem. A pool of Hydra runtimes can be used to accommodate any function invocation supported by Hydra.
@@
Virtualizing Runtimes allows one or more functions to share a single runtime. Sharing, however, can be achieved with different levels of isolation. Crucial (pons:2019, ) and Nightcore (jia:2021, ) handle multiple function invocations in a single runtime using different threads. Besides, Nightcore uses forking to spawn runtime processes, effectively eliminating support for such languages as Java. Photons (dukic:2020, ) also runs concurrent function invocations in a single runtime but relies on Java-specific bytecode transformations to ensure isolation (i.e., concurrent function invocations could not interact with other co-executing invocations). Boucher et al. (boucher:2018, ) rely on Rust’s memory safety guarantees to handle multiple invocations concurrently in isolation. Cloudflare Workers (cloudflare, ) rely on V8 memory isolates to run concurrent JavaScript functions in the same V8 runtime instance. Faasm (shillaker:2020, ; segarra:2025, ), Sledge (gadepalli:2020, ), and Gackstatter et al. (gackstatter:2022, ) proposed using WebAssembly (haas:2017, ) to execute multiple invocations in isolation. WebAssembly, however, still presents some challenges when running high-level languages such as Java, JavaScript, and Python (see section 2.2 for more details). As a result, runtimes such as Faasm need to compile a Python runtime such as CPython to WebAssembly in order to run Python functions. If concurrent invocations are to be handled, multiple runtimes will run concurrently in different WebAssembly execution environments, leading to resource redundancy.
@@
Hydra enforces isolation through memory sandboxes. However, compared to Cloudflare Workers, Hydra is not limited to JavaScript or WebAssembly functions. Hydra also supports applications that carry native dependencies through forked sandboxes and also supports sandbox snapshotting. Opposed to Faasm (which is centered on WebAssembly-based functions), Hydra relies on Truffle to support multiple language support, and therefore, concurrent invocations always share the underlying runtime. In addition, to further share the JIT compilation and profiling, Hydra shares the code caches among concurrent invocations of the same function.
@@
Profile and Code Cache Sharing reduce code profiling and compilation overheads. Profile sharing has been proposed by Arnold et al. (arnold:2005, ) and Ottoni et al. (ottoni:2021, ) as a way to re-use profiling information. By doing so, future invocations do not need to repeat all the code profiling steps and can benefit from the profiles gathered by previous invocations. A similar technique has also been used to share Just-In-Time compiled code (khrabrov:2022, ; xu:2018, ) among different invocations of the same application/function. Such techniques could be integrated into Hydra to share Truffle code caches. However, solutions such as JITServer (khrabrov:2022, ) and ShareJIT (xu:2018, ) are designed to optimize long-running applications and are not prepared to deal with the challenges inherent to fast-executing and lightweight serverless functions (carreira:2021, ).